Molecular iodine, I2(g), dissociates into iodine atoms at 625 K with a first-order rate constant of 0.271 s-1.

(a) What is the half-life for this reaction?

_______ s

(b) If you start with 0.051 M I2 at this temperature, how much will remain after 5.30 s assuming that the iodine atoms do not recombine to form I2?

_______ M
Answer #1

half-life = 0.693 / 0.271 = 2.5571 Sec

ln A = -kt + ln Ao

ln A = - 0.271 x 5.3 + ln 0.051

ln A = 1.21 x 10-2 M